# Affordance-based Robot Manipulation with Flow Matching

### v1

@@ 5 Conclusion
We have formulated a prompt tuning method for affordance learning and flow matching policy for robot manipulation. The core idea of prompt tuning is to maximally exploit the pretrained foundation model, and rapidly excavate the relevance of foundation and downstream affordance learning tasks. Our focus in this work is not to outperform state-of-the-art general robot manipulation research. Instead, we have systematically studied flow matching framework, which provides an alternative to diffusion policy. The results suggest forsaking the stochastic construction in favor of more directly learning the probability path, allowing for improved generation. We qualitatively and quantitatively experiment on the multi-task robot manipulation scenarios to prove the ease of training and evaluation for flow matching. We could further improve performance by incorporating the robot state as an extra input of flow matching, estimating closed-loop short horizon motions, and considering depth or pointcloud information for 3D orientation modeling. We leave these for future work.

### v2

@@ 5 Conclusion
We have formulated a prompt tuning method for affordance learning and flow matching policy for robot manipulation. The core idea of prompt tuning is to maximally exploit the pretrained foundation model, and rapidly excavate the relevance of foundation and downstream affordance learning tasks. Our focus in this work is not to outperform state-of-the-art general robot manipulation research. Instead, we have systematically studied flow matching framework, which provides an alternative to diffusion policy. The results suggest forsaking the stochastic construction in favor of learning the probability path more directly, allowing for improved generation and faster inference. We qualitatively and quantitatively experiment on the multi-task robot manipulation scenarios to prove the ease of training and evaluation for flow matching.
The 3D estimation performs decently with our method using only RGB images as input. But we also observe the 3D estimation could be highly attuned to camera angles. We could further improve performance by incorporating depth or pointcloud information for 3D orientation modeling. We leave these for future work.